# Mutual Information Estimation via Score-to-Fisher Bridge for Nonlinear Gaussian Noise Channels

### v1

@@ IV-C Log-domain trapezoid integration
Rather than applying the trapezoid rule directly on the non-uniform $t$-grid, we perform the change of variables $u=\log t$ to obtain uniform spacing in the log domain. Using the integral representation
$I(X;Y_{T})=\frac{1}{2}\int_{T}^{\infty}\!\left(\frac{n}{t}-J(Y_{t})\right)\,dt,$ (45)
and substituting $u=\log t$ (so $dt=e^{u}du$ and $t=e^{u}$), we obtain
@@
Since the $t$-grid (44) corresponds to uniform spacing in $u=\log t$ with step size $\Delta_{u}\equiv\log(t_{\max}/t_{\min})/(M-1)$, we apply the standard trapezoid rule (uniform spacing):
@@
where $\ell(u)\equiv n-e^{u}\,\widehat{J}(Y_{e^{u}})$ (equivalently $\ell(t)=n-t\,\widehat{J}(Y_{t})$), and $k_{T}$ is the smallest index with $t_{k_{T}}\geq T$. This log-domain integration offers several benefits over direct $t$-domain trapezoid integration: (1) uniform grid spacing reduces discretization error, (2) the transformation naturally handles a wide dynamic range in $t$, and (3) convexity bias is significantly reduced. In our experiments, this approach improved MI estimation accuracy by over an order of magnitude.
@@ V Nonlinear channel
where $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is a deterministic function. The mutual information equivalence presented in the following proposition is a key observation which enables the score-based estimation to be applied to general nonlinear channels.
@@ V-A Mutual information equivalence
Let $W\equiv f(X)$ be a measurable deterministic mapping, and consider $Y_{t}=W+Z_{t}$ with $Z_{t}\sim\mathcal{N}(\bm{0},t\bm{I}_{n})$ independent of $(X,W)$. Then the conditional law of $Y_{t}$ given $X$ depends on $X$ only through $W$. This means that $X\to W\to Y_{t}$ forms a Markov chain and $Y_{t}\perp X\mid W$.
@@ V-B Score-based estimation method for nonlinear channels
In the following discussin, we assume that the evalution of $f$ is computationally feasible. Namely, sampling from $p_{W}$ is feasible. Some parts of Algorithm 1 are replaced by the following steps:
  • Sample $W$ by $W=f(X)$, where $X\sim p_{X}$.
  • Estimate the Fisher information of $Y_{t}$ from the samples of $W$ by the score-based estimation method in Algorithm 1.
  • Use the estimated Fisher information to evaluate the integral form of MI.
  • Use the relationship $I(X;Y_{t})=I(W;Y_{t})$ to obtain the estimate of $I(X;Y_{t})$.
The estimation procedure is still a single-loop procedure, which does not require the access to the posterior $p_{X|Y_{t}}$.
@@ VI-A1 Experimental Setup
We implemented the DSM method with the following configuration: The score network is implemented as a fully connected neural network consisting of 3 hidden layers, each containing 128 units with SiLU (Swish) activation functions. The input and output dimensions match the problem dimension $n$, where we test with $n\in\{4,8,16\}$ to evaluate scalability. For the training configuration, we set the signal power to $P=1.0$ and construct a logarithmically spaced grid of $M=10$ noise variance values ranging from $t\in[P/200,200P]$, which covers four orders of magnitude. At each noise level $t$, we train a separate score network for 300 iterations using mini-batches of 4096 samples. The optimization is performed using the Adam optimizer with a learning rate of $10^{-3}$, and gradient clipping is applied with a maximum norm of 1.0 to ensure stable training.
@@ VI-A2 Validating the Fisher Information Estimate
Figure 1: Fisher information $J(Y_{t})$ versus noise variance $t$ for Gaussian input with $P=1$ and dimensions $n\in\{4,8,16\}$. Solid lines with circular markers show ground truth values $J(Y_{t})=n/(P+t)$, while dashed lines with triangular markers show DSM estimates $\hat{J}(Y_{t})$.
@@
VI-A3 Validating the Fisher-based MMSE Estimate
To validate our Fisher-based MMSE estimation, we compare three methods using the same experimental setup with $n=4$: (1) true MMSE (analytical ground truth), (2) Tweedie-based direct estimation using
$\widehat{\text{mmse}}_{\text{dir}}(t)\equiv\frac{1}{N}\sum_{i=1}^{N}\|\bm{x}_{i}-(\bm{y}_{i}+ts_{\bm{\theta}}(\bm{y}_{i}))\|^{2},$ (51)
and (3) our Fisher-based method using
$\widehat{\text{mmse}}_{\text{F}}(t)=nt-t^{2}\hat{J}(Y_{t}).$ (52)
The noise variance range is set to $t\in[P/200,50P]$ to avoid numerical instabilities at extremely high noise levels. The results are shown in Figure 2. The two estimates are in excellent agreement with the true MMSE in the wide range of noise variance.
Figure 2: Comparison of MMSE estimation methods for Gaussian input ($n=4$, $P=1$).
VI-A4 Mutual Information Estimation
We demonstrate the complete pipeline from the Score-to-Fisher bridge for mutual information estimation using the integral representation of MI. The experiment uses the same configuration as previous tests with $n=4$ and $P=1.0$, implementing log-domain trapezoid integration with tail correction. Figure 3 presents the estimated mutual information (per symbol) $\hat{I}(X;Y_{t})/n$ compared with the analytical ground truth $I(X;Y_{t})/n=(1/2)\log(1+P/t)$ for Gaussian input. The DSM-based estimates demonstrate excellent agreement with the theoretical values across the entire noise variance range.
@@ VI-A4 Mutual Information Estimation
VI-A5 Noise-conditional model
@@
For Gaussian input with $n=4$ and $P=1.0$, this approach achieves 6.61% median relative error across 12 evaluation points spanning $t\in[0.005,50]$. Figure 4 demonstrates reasonable agreement with analytical values across four orders of magnitude in noise variance. We can observe the tendency that the conditional DSM estimate slightly underestimates the MI at low noise variance.
@@
This unified framework requires only one neural network instead of 12 separate models, significantly reducing computational cost while eliminating the extreme outliers observed in per-$t$ training (Scheme A) and providing smooth interpolation across the entire noise range.
@@ VI-C Linear Gaussian channel case
$Y_{t}=AX+Z_{t},\qquad Z_{t}\sim\mathcal{N}(\mathbf{0},t\mathbf{I}_{n}),$ (53)
where $A\in\mathbb{R}^{n\times n}$ is a random orthogonal matrix and $X\sim\mathcal{N}(\mathbf{0},P\mathbf{I}_{n})$ with $P=1.0$, $n=4$.
@@
Figure 6: Mutual information estimation for linear Gaussian channel $Y_{t}=AX+Z_{t}$ ($n=4$, $P=1$). Circles: analytical ground truth; triangles: DSM estimates.
@@ VI-D Nonlinear channel case
$Y_{t}=\tanh(AX)+Z_{t},\quad Z_{t}\sim\mathcal{N}(\mathbf{0},t\mathbf{I}_{n}),$ (54)
where $A\in\mathbb{R}^{n\times n}$ is a random orthogonal matrix and $\tanh(\cdot)$ is applied elementwise. This model represents practical scenarios where signals undergo linear transformations before nonlinear processing.
@@ VI-D1 Experimental Setup
We use the same DSM configuration as in the linear case: $n=4$, $P=1.0$, with $M=12$ logarithmically spaced noise levels in $t\in[P/200,50P]$. The orthogonal matrix $A$ is generated via QR decomposition of a random Gaussian matrix. For each $t$, we train score networks for 400 iterations using 8192-sample mini-batches. Fisher information estimation uses 100,000 Monte Carlo samples per noise level.
@@ VI-D3 Results and Discussion
Figure 7: Mutual information estimation for composite nonlinear channel $Y_{t}=\tanh(AX)+Z_{t}$ ($n=4$, $P=1$). Circles: KDE-LOO baseline; triangles: DSM estimates.
@@
    Computational efficiency: While KDE-LOO requires $O(N^{2})$ distance computations per noise level, DSM achieves comparable accuracy with $O(N)$ score evaluations after one-time training.
@@
VI-E Internal validation via Fisher information
To further validate the theoretical foundations of our approach, we conduct an internal consistency check by estimating Fisher information through two independent routes within the proposed framework. For the same composite nonlinear channel $Y_{t}=\tanh(AX)+Z_{t}$, we compare:
  1. 1.
    Score-based route: $J_{\text{score}}(Y_{t})\equiv\mathbb{E}[\|s_{\bm{\theta}}(Y_{t})\|^{2}]$ from trained score networks
  2. 2.
    MMSE-based route: $J_{\text{mmse}}(Y_{t})\equiv{(nt-\text{mmse}_{W}(t))}/{t^{2}}$ from trained denoisers
Both routes should yield identical Fisher information if the theoretical framework is sound.
Figure 8 presents the comparison between the two estimation routes across the noise variance range. At the low noise variance, the MMSE-based route underestimates the Fisher information due to the difficulty of the learning of the optimal denoiser at low noise variance. However, at the mid to high noise variance, the two routes show excellent agreement, with both curves nearly overlapping throughout most of the range. This internal consistency check provides strong evidence for the correctness of our theoretical approach and implementation.
Figure 8: Internal validation via Fisher information estimation. Circles: score-based route; triangles: MMSE-based route.
This validation experiment is particularly valuable as it provides a self-contained verification of the method’s theoretical consistency, independent of external baselines or ground truth comparisons.
@@ VII Conclusion
We have presented a novel approach for mutual information estimation in nonlinear Gaussian noise channels via the Score-to-Fisher bridge. The method leverages DSM to estimate Fisher information, which is then converted to mutual information through the I-MMSE relation and log-domain integration with tail correction.
Experimental validation across linear Gaussian channels, composite nonlinear channels, and internal Fisher consistency checks demonstrates excellent agreement with analytical solutions and baseline methods. The approach achieves computational efficiency with $O(N)$ evaluations per noise level while maintaining high accuracy across wide dynamic ranges of the noise variance.
The Score-to-Fisher bridge methodology provides a practical and theoretically sound framework for mutual information estimation in realistic nonlinear Gaussian noise channel models where closed-form solutions are unavailable. Accurate estimation of mutual information plays a crucial role in assessing the fundamental limits of coded communication systems, and also serves as a valuable guideline for system design choices such as coding schemes, preprocessing strategies, and detection or decoding algorithms.

### v2

@@ IV-B Log-domain trapezoid integration
Rather than applying the trapezoid rule directly on the non-uniform $t$-grid, we perform the change of variables $u=\log t$ to obtain uniform spacing in the log domain. Using the integral representation (4), and substituting $u=\log t$ (so $dt=e^{u}du$ and $t=e^{u}$), we obtain
@@
Since the $t$-grid (42) corresponds to uniform spacing in $u=\log t$ with step size $\Delta_{u}\equiv\log(t_{\max}/t_{\min})/(M-1)$, we apply the standard trapezoid rule with uniform spacing:
@@
where $\ell(u)\equiv n-e^{u}\,\widehat{J}(Y_{e^{u}})$, and $k_{T}$ is the smallest index with $t_{k_{T}}\geq T$. This log-domain integration offers several benefits over direct $t$-domain trapezoid integration: (1) uniform grid spacing reduces discretization error, (2) the transformation naturally handles a wide dynamic range in $t$, and (3) convexity bias is significantly reduced. In our experiments, this approach improved MI estimation accuracy by over an order of magnitude.
@@ V Nonlinear channel
where $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is a deterministic function [^2] An extension to a general case $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ is straightforward but we assume $m=n$ for simplicity.. The mutual information equivalence presented in the following proposition is a key observation which enables the score-based estimation to be applied to general nonlinear channels.
@@ V-A Mutual information equivalence
Let $W\equiv f(X)$ be a measurable deterministic mapping, and consider $Y_{t}=W+Z_{t}$ with $Z_{t}\sim\mathcal{N}(\bm{0},t\bm{I}_{n})$ independent of $(X,W)$. Then the conditional law of $Y_{t}$ given $X$ depends on $X$ only through $W$. This means that $X\to W\to Y_{t}$ forms a Markov chain and equivalently $Y_{t}\perp X\mid W$.
@@ V-B Score-based estimation method for nonlinear channels
In the following discussin, we assume that the evalution of $f$ is computationally feasible. Namely, sampling from $p_{W}$ is feasible. Some parts of Algorithm 1 should be replaced by the following steps:
  • Sample $W$ by $W=f(X)$ where $X\sim p_{X}$.
  • Estimate the Fisher information of $Y_{t}$ from the samples of $W$ by the score-based estimation in Algorithm 1.
  • Use the estimated Fisher information to evaluate the Fisher integral form of MI.
  • Use the equality $I(X;Y_{t})=I(W;Y_{t})$ to obtain the estimate of $I(X;Y_{t})$.
The estimation procedure is still a single-loop procedure, which does not require the access to the posterior density $p_{X|Y_{t}}(\cdot|\cdot)$.
@@ VI-A1 Experimental Setup
We implemented the DSM learning for $p_{Y_{t}}$ with the following configuration: The score network is implemented as a fully connected neural network consisting of 3 hidden layers, each containing 128 units with SiLU (Swish) activation functions. The input and output dimensions match the problem dimension $n$, where we test with $n\in\{4,8,16\}$ to evaluate scalability. For the training configuration, we set the signal power to $P=1.0$ and construct a logarithmically spaced grid of $M=10$ noise variance values ranging from $t\in[P/200,200P]$, which covers four orders of magnitude. At each noise level $t$, we train a separate score network for 300 iterations using mini-batches of 4096 samples. The optimization is performed using the Adam optimizer with a learning rate of $10^{-3}$, and gradient clipping is applied with a maximum norm of 1.0 to ensure stable training.
@@ VI-A2 Validating the Fisher Information Estimate
Figure 1: Fisher information $J(Y_{t})$ versus noise variance $t$ for Gaussian input with $P=1$ and dimensions $n\in\{4,8,16\}$. Solid lines with circular markers show ground truth values $J(Y_{t})=n/(P+t)$, while dashed lines with triangular markers show estimated values $\hat{J}(Y_{t})$.
@@
VI-A3 Mutual Information Estimation
We demonstrate the complete pipeline from the Score-to-Fisher bridge for MI estimation using the integral representation of MI. The experiment uses the same configuration as previous tests with $n=4$ and $P=1.0$, implementing log-domain trapezoid integration with tail correction. Figure 2 presents the estimated mutual information (per symbol) $\hat{I}(X;Y_{t})/n$ compared with the analytical ground truth $I(X;Y_{t})/n=(1/2)\log(1+P/t)$ for Gaussian input. The DSM-based estimates demonstrate excellent agreement with the theoretical values across the entire noise variance range.
@@ VI-A3 Mutual Information Estimation
VI-A4 Noise-conditional model
@@
For Gaussian input with $n=4$ and $P=1.0$, this approach achieves 6.61% median relative error across 12 evaluation points spanning $t\in[0.005,50]$. Figure 3 demonstrates reasonable agreement with analytical values across four orders of magnitude in noise variance. We can observe the tendency that the noise-conditional DSM estimate slightly underestimates the MI at low noise variance. This result may be due to insufficient training iterations for a conditional model. However, this unified framework requires only one neural network instead of 12 separate models, significantly reducing computational cost.
@@ VI-C Linear Gaussian channel case
$Y_{t}={\bm{A}}X+Z_{t},\qquad Z_{t}\sim\mathcal{N}(\mathbf{0},t\mathbf{I}_{n}),$ (48)
where ${\bm{A}}\in\mathbb{R}^{n\times n}$ is a random orthogonal matrix and $X\sim\mathcal{N}(\mathbf{0},P\mathbf{I}_{n})$ with $P=1.0$, $n=4$.
@@
Figure 5: Mutual information estimation for linear Gaussian channel $Y_{t}={\bm{A}}X+Z_{t}$ ($n=4$, $P=1$). Circles: analytical ground truth; triangles: DSM estimates.
@@ VI-D Nonlinear channel case
$Y_{t}=\tanh({\bm{A}}X)+Z_{t},\quad Z_{t}\sim\mathcal{N}(\mathbf{0},t\mathbf{I}_{n}),$ (49)
where ${\bm{A}}\in\mathbb{R}^{n\times n}$ is a random orthogonal matrix and $\tanh(\cdot)$ is applied elementwise. This model represents practical scenarios where signals undergo linear transformations before nonlinear processing.
@@ VI-D1 Experimental Setup
We use the same DSM configuration as in the linear case: $n=4$, $P=1.0$, with $M=12$ logarithmically spaced noise levels in $t\in[P/200,50P]$. The orthogonal matrix $\bm{A}$} is generated via QR decomposition of a random Gaussian matrix. For each $t$, we train score networks for 400 iterations using 8192-sample mini-batches. Fisher information estimation uses 100,000 Monte Carlo samples per noise level.
@@ VI-D3 Results and Discussion
Figure 6: Mutual information estimation for composite nonlinear channel $Y_{t}=\tanh({\bm{A}}X)+Z_{t}$ ($n=4$, $P=1$). Circles: KDE-LOO baseline; triangles: DSM estimates.
@@
    Computational efficiency: While KDE-LOO requires $O(N^{2})$ distance computations per noise level, DSM achieves comparable accuracy with $O(N)$ score evaluations after one-time training where $N$ is the number of samples.
@@ VII Conclusion
We have presented a novel approach for mutual information estimation in nonlinear Gaussian noise channels via the Score-to-Fisher bridge approach. The method leverages DSM learning of score model for $p_{Y_{t}}$ to estimate Fisher information, which is then converted to mutual information via the Fisher integral representation.
Experimental validation across linear Gaussian channels, composite nonlinear channels, demonstrates excellent agreement with analytical solutions and baseline methods. The approach achieves computational efficiency with $O(N)$ evaluations per noise level while maintaining high accuracy across wide dynamic ranges of the noise variance. The curse of dimensionality can be avoided due to the efficiency of the score model evaluation.
The Score-to-Fisher bridge methodology provides a practical and theoretically sound framework for MI estimation in realistic nonlinear Gaussian noise channel models where closed-form solutions are unavailable. Accurate estimation of MI plays a crucial role in assessing the fundamental limits of coded communication systems, and also serves as a valuable guideline for system design choices such as coding schemes, preprocessing strategies, and detection or decoding algorithms.